# On the effects of radiation on mass transfer in binary stars

### v1

@@ 4.1 Relation to hydrostatic structure
An important aspect of our model is the connection between the hydrodynamical mass flow around L1 and the underlying hydrostatic donor envelope structure calculated using a stellar evolution code. Similarly to CP23, we notice that all the limiting cases of our model that we are able to solve, sets of equations (20) and (22), can be reformulated so that the Roche potential $\phi_{\rm R}$ becomes the independent variable instead of the coordinate $x$, allowing us to completely omit the coordinate $x$. This is what we do in practice in this paper. The mapping between $\phi$ and $r$ is based on the assumption that equipotentials can be related to spheres with the same equivalent volume. In this work, we use the higher order approximation derived by Jackson et al. (2017) to make the correspondence, see equation (67). We denote the radius corresponding to the inner boundary of our model at $\phi_{0}$, or also $x_{0}$, by $R_{0}$. The radius corresponding to the equipotential passing through the L1 point with $\phi_{1}$ is $R_{\text{L}}$.
@@
If the photosphere is below the L1 point, $R_{\rm d}<R_{\rm L}$, we have a Roche-lobe-underfilling donor and an optically thin regime. In this case, we take the inner boundary of our model to be the donor’s photosphere (unlike CP23) and we calculate the potential difference corresponding to $R_{0}=R_{\rm d}$ according to
@@
Finally, let us also explain how we calculate the Eddington factor $\Gamma_{\rm Edd}$. The usage of the von Zeipel theorem does not change the value of the Eddington factor. Thus, in the optically thick regime in the donor’s interior we calculate it according to (e.g. Sanyal et al., 2015)
@@
where $\bar{P}_{\rm rad}$ is the hydrostatic radiative pressure and ¯$\bar{\nabla}_{T}$ is the actual hydrostatic temperature gradient. In the optically thin regime outside the donor we use a simple model, where we assume $\Gamma_{\rm Edd}$ to be constant and take the value at the photosphere.
@@ 4.2 Analytic estimates
The sets of equations (20) or (22) cannot be analytically solved for a general EOS, but we can still analytically estimate the quantity of main interest, $-\dot{M}_{\rm d}=Q\rho v$ (equation 13). We need to know the underlying hydrostatic structure of the donor from a 1D stellar evolution code, such as the hydrostatic stellar quantities $\bar{\rho}$, $\bar{P}_{\rm gas}$ as a function of stellar radius $r$ or alternatively the stellar gravitational potential $\bar{\phi}$. We also need to know the gas cross-section $Q$. We need to evaluate $\dot{M}_{\rm d}$ at one point, and the easiest choice is the critical point located where the right-hand side of equation (20b) vanishes. If the Eddington limit is not reached, then the critical point occurs at the Roche-lobe radius $R_{\rm L}$ that corresponds to $\bar{\phi}(R_{\rm L})=\phi_{1}$, where $\phi_{1}$ is the Roche potential at L1. Combining equations (20a) and (20b) we see that the critical velocity $v_{\rm crit}\equiv v(x_{\rm crit})$ is given by $v_{\rm crit}^{2}=c_{T,\rm{gas}}^{2}=P_{\rm gas}/\rho$. Now, we need to express the hydrodynamic values $P_{\rm gas}$, $\rho$, which we a priori do not know, as functions of the hydrostatic values $\bar{P}_{\rm gas}$, $\bar{\rho}$, which we know from a stellar evolution code. Considering the simple case of an isothermal gas (see Sec. 5.1 in CP23, or also Sec. 3.1.3 in Lamers & Cassinelli 1999) we can approximately write $P_{\rm gas}=\bar{P}_{\text{gas}}/\sqrt{\text{e}}$ and $\rho=\bar{\rho}/\sqrt{\rm e}$ at the critical point.
@@
where the centers of the two stars lie on the $x$ axis and the orbital plane is the $xy$ plane (CP23). The cross-section $Q$ can then be approximated as
@@
If we assume $\Gamma_{\rm Edd}$ to be roughly constant in the $yz$-plane then the potential is effectively lower by a factor of $(1-\Gamma_{\rm Edd})$ and it is equivalent to assume $B\to(1-\Gamma_{\rm Edd})B$ and $C\to(1-\Gamma_{\rm Edd})C$. In such a case we can write (see the derivation of $Q_{\rho}$ in CP23)
@@
$-\dot{M}_{\rm d}\approx\dot{M}_{\rm an,thick}\equiv\left.\frac{1}{1-\Gamma_{\rm Edd}}\frac{1}{\sqrt{e}}\frac{2\pi}{\sqrt{BC}}\frac{\bar{P}_{\rm gas}^{3/2}}{\bar{\rho}^{1/2}}\right|_{r=R_{\rm L}}.$ (32)
@@ 4.3.1 Isothermal gas
The only difference between this isothermal limit and the isothermal limit in CP23 is that the solution here includes the momentum exchange between gas and radiation. Therefore, the potential difference the gas has to cover between $x_{0}$ and $x_{1}$ (L1) is multiplied by a factor of $(1-\Gamma_{\rm Edd})$ and thus somewhat reduced: $(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}\to(1-\Gamma_{\rm Edd})(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}$. The isothermal case in CP23 is equivalent to the optically thin MT rate of Ritter (1988), or its improved version of Jackson et al. (2017), depending on how we choose to approximate the difference between the potentials at the L1 point $\phi_{1}$ and the photosphere $\phi_{\rm ph}$. Thus, the MT rate corresponding to the isothermal case with radiation is
@@ 4.3.2 Ideal gas
As before, compared to the ideal gas limit of CP23, the only difference is that the current ideal gas case includes momentum exchange between gas and radiation. Consequently, the potential difference is reduced: $(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}\to(1-\Gamma_{\rm Edd})(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}$. Therefore, we do not discuss this case further in this paper and refer the reader to CP23 for more details.
@@ 4.4 Numerical solution for a realistic equation of state
For a general EOS, we solve the set of hydrodynamic equations (20) or (22) using the same framework we developed in CP23 based on a relaxation method of Press et al. (2007, Chapter 18). The realistic EOS is adopted from the MESA eos module (Paxton et al., 2011, 2013, 2015, 2018, 2019). Depending on the metallicity $Z$ and the position in the $\rho T$ plane, one of the following components is used: OPAL/SCVH (Rogers & Nayfonov, 2002; Saumon et al., 1995), Free EOS (Irwin, 2008), HELM (Timmes & Swesty, 2000), PC (Potekhin & Chabrier, 2010), Skye (Jermyn et al., 2021), or CMS (Chabrier et al., 2019). Following the tutorial by Timmes (2021), we construct a table with all relevant thermodynamic variables. We interpolate in the table using bilinear interpolation (Press et al., 2007, Chapter 3). We refer the reader to our original work (CP23) for more details.
In what follows, we compare numerical solutions in the small-$v$ case (Section 3.2, equations 20), the adiabatic case (Section 3.3, equations 22) and our original model in CP23. First, in Section 4.4.1 we use the same low metallicity donor $30\,\rm{M}_{\odot}$ as in CP23 that is adopted from Marchant et al. (2021). Our results will justify the usage of the analytical estimate in the optically thick regime (equation 32). Next, in Section 4.4.2 we illustrate the differences between our present approach and the standard approach on a $40\,\rm{M}_{\odot}$ low-metallicity donor in a supergiant stage. This donor is ideal for this purpose because it achieves high values of the Eddington ratio in the surface layers ($\Gamma_{\rm Edd}>0.5$), but does not cross into the super-Eddington regime. For completeness, we present in Appendix D comparison for a $1\,\rm{M}_{\odot}$ red giant showing that in this case radiation has little effect on $\dot{M}$, as expected.
Finally, we introduce the notation for the previous MT rate prescriptions. For Roche-lobe-underfilling donors, $R_{\rm d}\leq R_{\rm L}$, we will use the optically thin model of Jackson et al. (2017), which is based on Ritter (1988). We will refer to this MT rate by $\dot{M}_{\rm J}$ and the expression can be found in equation (66). For a Roche-lobe-overflowing donor, $R_{\rm d}\geq R_{\rm L}$, we will use the optically thick model of Kolb & Ritter (1990). We will refer to this MT rate by $\dot{M}_{\rm KR}$ and the expression is in equation (68). In some cases, we will also use the prescription introduced by Marchant et al. (2021), $\dot{M}_{\rm M}$.
@@
We used a $30\,\rm{M}_{\odot}$ low-metallicity donor undergoing intensive thermal time-scale MT to a black hole, as was investigated by Marchant et al. (2021). The donor’s initial metallicity is $Z=\rm{Z}_{\odot}/10$, where $\rm{Z}_{\odot}=0.0142$ (Asplund et al., 2009) and where the relative metal mass fractions are from Grevesse & Sauval (1998). We take a model with a $7.5\,\rm{M}_{\odot}$ black hole and an initial period of $31.6$ d. We recalculated the evolution using the full MT prescription $\dot{M}_{\rm{M}}$ and the same MESA version as in Marchant et al. (2021).
Figure 2: Comparison of density $\rho$, temperature $T$, and velocity $v$ profiles for a $30\rm{M}_{\odot}$ low-metallicity star undergoing thermal time-scale MT (CP23), small-$v$ case (Section 3.2), and adiabatic case (Section 3.3). We also show the underlying MESA hydrostatic density and temperature profiles and the value of convective velocity $v_{\rm conv}$. The profiles are shown as functions of radius $r$ going from the inner boundary at $R_{0}$ (that corresponds to $x_{0}$), to the outer boundary at $R_{\rm L}$ (the L1 point at $x_{1}$). The critical velocity $v_{\rm crit}$, indicated by dotted lines, is reached at the critical point, L1. With the top $x$ axis we indicate the distance to the critical point in units of pressure scale heights $\Delta N_{H_{P}}$.
As in CP23, we first examine the binary near the end of the thermal time-scale MT phase at an age of $6.79\,\rm{Myr}$, with a surface Eddington factor $\Gamma_{\rm Edd}=0.78$, donor mass $M_{\rm{d}}=14.0\,\rm{M}_{\odot}$, core helium abundance $0.36$, relative radius excess $\delta R_{\text{d}}=3.18\times 10^{-2}$, and mass ratio $q=1.85$. Here, $\delta R_{\rm d}\equiv\Delta R_{\rm d}/R_{\rm L}$, where $\Delta R_{\rm d}\equiv R_{\rm d}-R_{\rm L}$ is the radius excess, and $q\equiv M_{\rm d}/M_{\rm a}$ is the donor-to-accretor (black hole) mass ratio.
@@
Figure 3: Comparison of the dependencies of the donor’s MT rate $-\dot{M}_{\rm d}$ on $R_{0}$ for the same cases and donor star as in Fig. 2. We also show the analytical estimate $\dot{M}_{\rm an,thick}$ (equation 32). The outer fixed boundary at $R_{L}$ is indicated as well as the distance between $R_{0}$ and $R_{\rm{L}}$ in units of pressure scale heights $\Delta N_{H_{P}}$.
In Fig. 3 we also see that the MT rates of CP23 and in the adiabatic case follow a similar trend, but the CP23 model gives a lower value of the MT rate, which we discuss in more detail in Section 4.5. The similarity of the trends is caused by the fact that the CP23 model is also adiabatic (in the sense that there is no energy exchange between gas and radiation), but it also includes the averaging of physical quantities over the L1 plane. Therefore, as the gas approaches L1, it cools down and its cross-section decreases, its potential energy decreases, and the difference is used to heat up the gas, as we can also see in the middle panel of Fig. 2.
In Fig. 4, we compare different MT rates throughout the evolution of the $30\,\rm{M}_{\odot}$ donor. First, we compare the small-$v$ case with the analytical expression $\dot{M}_{\rm an,thick}$ (equation 32). We then compare the MT prescription of Kolb & Ritter (1990), $\dot{M}_{\rm KR}$, with $\dot{M}_{\rm an,thick}$, followed by a comparison of the MT rate $\dot{M}_{\rm M}$ from Marchant et al. (2021) with the modified analytical expression
$-\dot{M}_{\rm d}\approx\dot{M}_{\rm an,thick,tot}\equiv\left.\frac{1}{\sqrt{e}}\frac{2\pi}{\sqrt{BC}}\frac{\bar{P}^{3/2}}{\bar{\rho}^{1/2}}\right|_{r=R_{\rm L}},$ (40)
@@
In the second panel, $\dot{M}_{\rm an,thick}$ agrees with $\dot{M}_{\rm KR}$ to within $\sim$20%, except at the final point, where $\dot{M}_{\rm an,thick}$ is approximately twice as large. In the third panel, $\dot{M}_{\rm an,thick,tot}$ matches $\dot{M}_{\rm M}$ within $\sim$10%, further justifying its use. Finally, the bottom panel shows that $\dot{M}_{\rm an,thick}$ is consistently about twice as low as $\dot{M}_{\rm an,thick,tot}$, which would impact binary evolution if used self-consistently instead of $\dot{M}_{\rm an,thick,tot}$ ($\dot{M}_{\rm M}$). The effects would resemble those of evolving the binary with $\dot{M}_{\rm KR}$ (see the second panel of Fig. 4). For details, refer to Marchant et al. (2021), or see CP23, where we previously illustrated the impact of reducing the MT rate by a factor of a few. Therefore, we do not repeat this analysis here but discuss the implications in more detail in Section 6.
Figure 4: Comparison of different MT prescriptions throughout the evolution of originally $30\rm{M}_{\odot}$ low-metallicity donor undergoing thermal time-scale MT (Section 4.4.1, Marchant et al. 2021). First panel: the small-$v$ case (Section 3.2) for three different values of the distances between the inner and outer boundary in units of pressure scale heights $\Delta N_{H_{P}}$ vs. the analytical estimate $\dot{M}_{\rm an,thick}$ (equation 32). Second panel: the MT prescription introduced by Kolb & Ritter (1990), $\dot{M}_{\rm KR}$ vs. the analytical estimate $\dot{M}_{\rm an,thick}$. Third panel: the MT prescription introduced by Marchant et al. (2021), $\dot{M}_{\rm M}$, vs. the analytical estimate $\dot{M}_{\rm an,thick,tot}$ (equation 40). Fourth panel: Comparison of the two analytical prescriptions $\dot{M}_{\rm an,thick,tot}$ vs. $\dot{M}_{\rm an,thick}$, throughout the evolution.
@@
In Fig. 5 we show the comparison of our new analytical estimates $\dot{M}_{\rm an,thin}$ (equation 36) and $\dot{M}_{\rm an,thick}$ (equation 32) with the standard model of optically-thin MT (Jackson et al., 2017, $\dot{M}_{\rm J}$, equation 66) and optically-thick MT (Kolb & Ritter, 1990, $\dot{M}_{\rm KR}$, equation 68). Here, we measure the absolute radius excess $\Delta R_{\rm d}$ in the units of the photosphere pressure scale height $H_{P,\rm ph}$ defined as
@@
where $P_{\rm ph}$ is the photosphere pressure. We see that in the optically thin regime $\dot{M}_{\rm an,thin}$ is greater than $\dot{M}_{\rm J}$ due to the momentum boost from radiation that is included in our expression, and the relative difference is growing exponentially. In contrast, $\dot{M}_{\rm an,thick}$ is similar to $\dot{M}_{\rm KR}$ because the model of Kolb & Ritter (1990) integrates the total pressure (including $P_{\rm rad}$) of the overflowing layers and therefore includes much of the radiation boost. We also confirm that $\dot{M}_{\rm an,thin}\approx\dot{M}_{\rm J}$ and $\dot{M}_{\rm an,thick}\approx\dot{M}_{\rm KR}$ for a gas pressure dominated envelope ($P_{\rm gas}\gg P_{\rm rad}$) on a solar-like donor on the red giant branch, see Appendix D.
Figure 5: Comparison of our analytical MT prescription $\dot{M}_{\rm an,thin}$ and $\dot{M}_{\rm an,thick}$ with the optically thin prescription introduced by Jackson et al. (2017), $\dot{M}_{\rm J},$ and the optically thick prescription introduced by Kolb & Ritter (1990), $\dot{M}_{\rm KR}$. We also show the Eddington factors $\Gamma_{\rm tot}=L/L_{\text{Edd}}$ and $\Gamma_{\rm Edd}$ (equation 28). The case shown is for an isolated originally $40\rm{M}_{\odot}$ low-metallicity star in the supergiant stage.

### v2

@@ 4.1 Relation to hydrostatic structure
An important aspect of our model is the connection between the hydrodynamical mass flow around L1 and the underlying hydrostatic donor envelope structure calculated using a stellar evolution code. Similarly to cehula2023, we notice that all the limiting cases of our model that we are able to solve, sets of equations (20) and (22), can be reformulated so that the Roche potential $\phi_{\rm R}$ becomes the independent variable instead of the coordinate $x$, allowing us to completely omit the coordinate $x$. This is what we do in practice in this paper. The mapping between $\phi$ and $r$ is based on the assumption that equipotentials can be related to spheres with the same equivalent volume. In this work, we use the higher order approximation derived by jackson2017 to make the correspondence, see equation (67). We denote the radius corresponding to the inner boundary of our model at $\phi_{0}$, or also $x_{0}$, by $R_{0}$. The radius corresponding to the equipotential passing through the L1 point with $\phi_{1}$ is $R_{\text{L}}$.
@@
If the photosphere is below the L1 point, $R_{\rm d}<R_{\rm L}$, we have a Roche-lobe-underfilling donor and an optically thin regime. In this case, we take the inner boundary of our model to be the donor’s photosphere (unlike cehula2023) and we calculate the potential difference corresponding to $R_{0}=R_{\rm d}$ according to
@@
Finally, let us also explain how we calculate the Eddington factor $\Gamma_{\rm Edd}$. The usage of the von Zeipel theorem does not change the value of the Eddington factor. Thus, in the optically thick regime in the donor’s interior we calculate it according to (e.g. sanyal2015)
@@
where $\bar{P}_{\rm rad}$ is the hydrostatic radiative pressure and $\bar{\nabla}_{T}$ is the actual hydrostatic temperature gradient. In the optically thin regime outside the donor we use a simple model, where we assume $\Gamma_{\rm Edd}$ to be constant and take the value at the photosphere.
@@ 4.2 Analytic estimates
The sets of equations (20) or (22) cannot be analytically solved for a general EOS, but we can still analytically estimate the quantity of main interest, $\dot{M}_{\rm d}=Q\rho v$ (equation 13). We need to know the underlying hydrostatic structure of the donor from a 1D stellar evolution code, such as the hydrostatic stellar quantities $\bar{\rho}$, $\bar{P}_{\rm gas}$ as a function of stellar radius $r$ or alternatively the stellar gravitational potential $\bar{\phi}$. We also need to know the gas cross-section $Q$. We need to evaluate $\dot{M}_{\rm d}$ at one point, and the easiest choice is the critical point located where the right-hand side of equation (20b) vanishes. If the Eddington limit is not reached, then the critical point occurs at the Roche-lobe radius $R_{\rm L}$ that corresponds to $\bar{\phi}(R_{\rm L})=\phi_{1}$, where $\phi_{1}$ is the Roche potential at L1. Combining equations (20a) and (20b) we see that the critical velocity $v_{\rm crit}\equiv v(x_{\rm crit})$ is given by $v_{\rm crit}^{2}=c_{T,\rm{gas}}^{2}=P_{\rm gas}/\rho$. Now, we need to express the hydrodynamic values $P_{\rm gas}$, $\rho$, which we a priori do not know, as functions of the hydrostatic values $\bar{P}_{\rm gas}$, $\bar{\rho}$, which we know from a stellar evolution code. Considering the simple case of an isothermal gas (see Sec. 5.1 in cehula2023, or also Sec. 3.1.3 in lamers1999) we can approximately write $P_{\rm gas}=\bar{P}_{\text{gas}}/\sqrt{\text{e}}$ and $\rho=\bar{\rho}/\sqrt{\rm e}$ at the critical point.
@@
where the centers of the two stars lie on the $x$ axis and the orbital plane is the $xy$ plane (cehula2023). The cross-section $Q$ can then be approximated as
@@
If we assume $\Gamma_{\rm Edd}$ to be roughly constant in the $yz$-plane then the potential is effectively lower by a factor of $(1-\Gamma_{\rm Edd})$ and it is equivalent to assume $B\to(1-\Gamma_{\rm Edd})B$ and $C\to(1-\Gamma_{\rm Edd})C$. In such a case we can write (see the derivation of $Q_{\rho}$ in cehula2023)
@@
$\dot{M}_{\rm d}\approx\dot{M}_{\rm an,thick}\equiv\left.\frac{1}{1-\Gamma_{\rm Edd}}\frac{1}{\sqrt{e}}\frac{2\pi}{\sqrt{BC}}\frac{\bar{P}_{\rm gas}^{3/2}}{\bar{\rho}^{1/2}}\right|_{r=R_{\rm L}}.$ (32)
@@ 4.3.1 Isothermal gas
The only difference between this isothermal limit and the isothermal limit in cehula2023 is that the solution here includes the momentum exchange between gas and radiation. Therefore, the potential difference the gas has to cover between $x_{0}$ and $x_{1}$ (L1) is multiplied by a factor of $(1-\Gamma_{\rm Edd})$ and thus somewhat reduced: $(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}\to(1-\Gamma_{\rm Edd})(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}$. The isothermal case in cehula2023 is equivalent to the optically thin MT rate of ritter1988, or its improved version of jackson2017, depending on how we choose to approximate the difference between the potentials at the L1 point $\phi_{1}$ and the photosphere $\phi_{\rm ph}$. Thus, the MT rate corresponding to the isothermal case with radiation is
@@ 4.3.2 Ideal gas
As before, compared to the ideal gas limit of cehula2023, the only difference is that the current ideal gas case includes momentum exchange between gas and radiation. Consequently, the potential difference is reduced: $(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}\to(1-\Gamma_{\rm Edd})(\phi_{1}-\phi_{0})/c_{T,\rm gas}^{2}$. Therefore, we do not discuss this case further in this paper and refer the reader to cehula2023 for more details.
@@ 4.4 Numerical solution for a realistic equation of state
For a general EOS, we solve the set of hydrodynamic equations (20) or (22) using the same framework we developed in cehula2023 based on a relaxation method of press2007. The realistic EOS is adopted from the MESA eos module (paxton2011; paxton2013; paxton2015; paxton2018; paxton2019). Depending on the metallicity $Z$ and the position in the $\rho T$ plane, one of the following components is used: OPAL/SCVH (rogers2002; saumon1995), Free EOS (irwin2008), HELM (timmes2000), PC (potekhin2010), Skye (jermyn2021), or CMS (chabrier2019). Following the tutorial by timmes2021, we construct a table with all relevant thermodynamic variables. We interpolate in the table using bilinear interpolation (press2007, Chapter 3). We refer the reader to our original work (cehula2023) for more details.
In what follows, we compare numerical solutions in the small-$v$ case (Section 3.2, equations 20), the adiabatic case (Section 3.3, equations 22) and our original model in cehula2023. First, in Section 4.4.1 we use the same low metallicity donor $30\,\rm{M}_{\odot}$ as in cehula2023 that is adopted from marchant2021. Our results will justify the usage of the analytical estimate in the optically thick regime (equation 32). Next, in Section 4.4.2 we illustrate the differences between our present approach and the standard approach on a $40\,\rm{M}_{\odot}$ low-metallicity donor in a supergiant stage. This donor is ideal for this purpose because it achieves high values of the Eddington ratio in the surface layers ($\Gamma_{\rm Edd}>0.5$), but does not cross into the super-Eddington regime. For completeness, we present in Appendix D comparison for a $1\,\rm{M}_{\odot}$ red giant showing that in this case radiation has little effect on $\dot{M}$, as expected.
Finally, we introduce the notation for the previous MT rate prescriptions. For Roche-lobe-underfilling donors, $R_{\rm d}\leq R_{\rm L}$, we will use the optically thin model of jackson2017, which is based on ritter1988. We will refer to this MT rate by $\dot{M}_{\rm J}$ and the expression can be found in equation (66). For a Roche-lobe-overflowing donor, $R_{\rm d}\geq R_{\rm L}$, we will use the optically thick model of kolb1990. We will refer to this MT rate by $\dot{M}_{\rm KR}$ and the expression is in equation (68). In some cases, we will also use the prescription introduced by marchant2021, $\dot{M}_{\rm M}$.
@@
We used a $30\,\rm{M}_{\odot}$ low-metallicity donor undergoing intensive thermal time-scale MT to a black hole, as was investigated by marchant2021. The donor’s initial metallicity is $Z=\rm{Z}_{\odot}/10$, where $\rm{Z}_{\odot}=0.0142$ (asplund2009) and where the relative metal mass fractions are from grevesse1998. We take a model with a $7.5\,\rm{M}_{\odot}$ black hole and an initial period of $31.6$ d. We recalculated the evolution using the full MT prescription $\dot{M}_{\rm{M}}$ and the same MESA version as in marchant2021.
Figure 2: Comparison of density $\rho$, temperature $T$, and velocity $v$ profiles for a $30\rm{M}_{\odot}$ low-metallicity star undergoing thermal time-scale MT (cehula2023), small-$v$ case (Section 3.2), and adiabatic case (Section 3.3). We also show the underlying MESA hydrostatic density and temperature profiles and the value of convective velocity $v_{\rm conv}$. The profiles are shown as functions of radius $r$ going from the inner boundary at $R_{0}$ (that corresponds to $x_{0}$), to the outer boundary at $R_{\rm L}$ (the L1 point at $x_{1}$). The critical velocity $v_{\rm crit}$, indicated by dotted lines, is reached at the critical point, L1. With the top $x$ axis we indicate the distance to the critical point in units of pressure scale heights $\Delta N_{H_{P}}$.
As in cehula2023, we first examine the binary near the end of the thermal time-scale MT phase at an age of $6.79\,\rm{Myr}$, with a surface Eddington factor $\Gamma_{\rm Edd}=0.78$, donor mass $M_{\rm{d}}=14.0\,\rm{M}_{\odot}$, core helium abundance $0.36$, relative radius excess $\delta R_{\text{d}}=3.18\times 10^{-2}$, and mass ratio $q=1.85$. Here, $\delta R_{\rm d}\equiv\Delta R_{\rm d}/R_{\rm L}$, where $\Delta R_{\rm d}\equiv R_{\rm d}-R_{\rm L}$ is the radius excess, and $q\equiv M_{\rm d}/M_{\rm a}$ is the donor-to-accretor (black hole) mass ratio.
@@
Figure 3: Comparison of the dependencies of the donor’s MT rate $\dot{M}_{\rm d}$ on $R_{0}$ for the same cases and donor star as in Fig. 2. We also show the analytical estimate $\dot{M}_{\rm an,thick}$ (equation 32). The outer fixed boundary at $R_{L}$ is indicated as well as the distance between $R_{0}$ and $R_{\rm{L}}$ in units of pressure scale heights $\Delta N_{H_{P}}$.
In Fig. 3 we also see that the MT rates of cehula2023 and in the adiabatic case follow a similar trend, but the cehula2023 model gives a lower value of the MT rate, which we discuss in more detail in Section 4.5. The similarity of the trends is caused by the fact that the cehula2023 model is also adiabatic (in the sense that there is no energy exchange between gas and radiation), but it also includes the averaging of physical quantities over the L1 plane. Therefore, as the gas approaches L1, it cools down and its cross-section decreases, its potential energy decreases, and the difference is used to heat up the gas, as we can also see in the middle panel of Fig. 2.
In Fig. 4, we compare different MT rates throughout the evolution of the $30\,\rm{M}_{\odot}$ donor. First, we compare the small-$v$ case with the analytical expression $\dot{M}_{\rm an,thick}$ (equation 32). We then compare the MT prescription of kolb1990, $\dot{M}_{\rm KR}$, with $\dot{M}_{\rm an,thick}$, followed by a comparison of the MT rate $\dot{M}_{\rm M}$ from marchant2021 with the modified analytical expression
$\dot{M}_{\rm d}\approx\dot{M}_{\rm an,thick,tot}\equiv\left.\frac{1}{\sqrt{e}}\frac{2\pi}{\sqrt{BC}}\frac{\bar{P}^{3/2}}{\bar{\rho}^{1/2}}\right|_{r=R_{\rm L}},$ (40)
@@
In the second panel, $\dot{M}_{\rm an,thick}$ agrees with $\dot{M}_{\rm KR}$ to within $\sim$20%, except at the final point, where $\dot{M}_{\rm an,thick}$ is approximately twice as large. In the third panel, $\dot{M}_{\rm an,thick,tot}$ matches $\dot{M}_{\rm M}$ within $\sim$10%, further justifying its use. Finally, the bottom panel shows that $\dot{M}_{\rm an,thick}$ is consistently about twice as low as $\dot{M}_{\rm an,thick,tot}$, which would impact binary evolution if used self-consistently instead of $\dot{M}_{\rm an,thick,tot}$ ($\dot{M}_{\rm M}$). The effects would resemble those of evolving the binary with $\dot{M}_{\rm KR}$ (see the second panel of Fig. 4). For details, refer to marchant2021, or see cehula2023, where we previously illustrated the impact of reducing the MT rate by a factor of a few. Therefore, we do not repeat this analysis here but discuss the implications in more detail in Section 6.
Figure 4: Comparison of different MT prescriptions throughout the evolution of originally $30\rm{M}_{\odot}$ low-metallicity donor undergoing thermal time-scale MT (Section 4.4.1, marchant2021). First panel: the small-$v$ case (Section 3.2) for three different values of the distances between the inner and outer boundary in units of pressure scale heights $\Delta N_{H_{P}}$ vs. the analytical estimate $\dot{M}_{\rm an,thick}$ (equation 32). Second panel: the MT prescription introduced by kolb1990, $\dot{M}_{\rm KR}$ vs. the analytical estimate $\dot{M}_{\rm an,thick}$. Third panel: the MT prescription introduced by marchant2021, $\dot{M}_{\rm M}$, vs. the analytical estimate $\dot{M}_{\rm an,thick,tot}$ (equation 40). Fourth panel: Comparison of the two analytical prescriptions $\dot{M}_{\rm an,thick,tot}$ vs. $\dot{M}_{\rm an,thick}$, throughout the evolution.
@@
In Fig. 5 we show the comparison of our new analytical estimates $\dot{M}_{\rm an,thin}$ (equation 36) and $\dot{M}_{\rm an,thick}$ (equation 32) with the standard model of optically-thin MT (jackson2017, $\dot{M}_{\rm J}$, equation 66) and optically-thick MT (kolb1990, $\dot{M}_{\rm KR}$, equation 68). Here, we measure the absolute radius excess $\Delta R_{\rm d}$ in the units of the photosphere pressure scale height $H_{P,\rm ph}$ defined as
@@
where $P_{\rm ph}$ is the photosphere pressure. We see that in the optically thin regime $\dot{M}_{\rm an,thin}$ is greater than $\dot{M}_{\rm J}$ due to the momentum boost from radiation that is included in our expression, and the relative difference is growing exponentially. In contrast, $\dot{M}_{\rm an,thick}$ is similar to $\dot{M}_{\rm KR}$ because the model of kolb1990 integrates the total pressure (including $P_{\rm rad}$) of the overflowing layers and therefore includes much of the radiation boost. We also confirm that $\dot{M}_{\rm an,thin}\approx\dot{M}_{\rm J}$ and $\dot{M}_{\rm an,thick}\approx\dot{M}_{\rm KR}$ for a gas pressure dominated envelope ($P_{\rm gas}\gg P_{\rm rad}$) on a solar-like donor on the red giant branch, see Appendix D.
Figure 5: Comparison of our analytical MT prescription $\dot{M}_{\rm an,thin}$ and $\dot{M}_{\rm an,thick}$ with the optically thin prescription introduced by jackson2017, $\dot{M}_{\rm J},$ and the optically thick prescription introduced by kolb1990, $\dot{M}_{\rm KR}$. We also show the Eddington factors $\Gamma_{\rm tot}=L/L_{\text{Edd}}$ and $\Gamma_{\rm Edd}$ (equation 28). The case shown is for an isolated originally $40\rm{M}_{\odot}$ low-metallicity star in the supergiant stage.